i dont know the answer to 26m,104m to the second power i have to find the LCM
Answer by stanbon(75887)

i dont know the answer to 26m,104m to the second power i have to find the LCM
--------------------
Write the "prime-factor form" for each term:
26m = 2*13*m
104m= 2^3*13*m^2
-----------------------
The LCM must contain each different factor in its highest power:
LCM = 2^3*13*m^2 = 104m^2
======================
cheers,
Stan H.